### v1

# MoEMeta: Mixture-of-Experts Meta Learning for Few-Shot Relational Learning

Abstract
Few-shot knowledge graph relational learning seeks to perform reasoning over relations given only a limited number of training examples. While existing approaches largely adopt a meta-learning framework for enabling fast adaptation to new relations, they suffer from two key pitfalls. First, they learn relation meta-knowledge in isolation, failing to capture common relational patterns shared across tasks. Second, they struggle to effectively incorporate local, task-specific contexts crucial for rapid adaptation. To address these limitations, we propose MoEMeta, a novel meta-learning framework that disentangles globally shared knowledge from task-specific contexts to enable both effective generalization and rapid adaptation. MoEMeta introduces two key innovations: (i) a mixture-of-experts (MoE) model that learns globally shared relational prototypes to enhance generalization, and (ii) a task-tailored adaptation mechanism that captures local contexts for fast task-specific adaptation. By balancing global generalization with local adaptability, MoEMeta significantly advances few-shot relational learning. Extensive experiments and analyses on three KG benchmarks demonstrate that MoEMeta consistently outperforms existing baselines, achieving state-of-the-art performance. [^1] Code is available at: https://github.com/alexhw15/MoEMeta.git.
1 Introduction
Knowledge graphs (KGs) encode real-world knowledge in the form of factual triplets $(h,r,t)$, where each triplet represents a relationship $r$ between a head entity $h$ and a tail entity $t$ (Bollacker et al., 2008; Suchanek et al., 2007; Vrandečić and Krötzsch, 2014). KGs are crucial for powering various intelligent applications, such as question answering (Wu et al., 2016), Web search (Fang et al., 2017), and advanced recommendation systems (Cao et al., 2019). Despite their widespread adoption, KGs are inherently incomplete. Although numerous KG completion methods  (Bordes et al., 2013; Yang et al., 2015; Trouillon et al., 2016) have been proposed to infer missing facts based on observed triplets, their effectiveness is severely hindered by data sparsity, especially for long-tail or emerging relations with only a handful of known training triplets.
In response, few-shot relational learning (FSRL) has emerged as a promising direction for reasoning over few-shot relations with limited training triplets, enabling generalization to novel relations in data-scarce scenarios. Most existing FSRL methods  (Chen et al., 2019; Jiang et al., 2021; Niu et al., 2021; Wu et al., 2023) build upon model-agnostic meta-learning (MAML) (Finn et al., 2017), a gradient-based meta-learning framework that learns a set of global parameters serving as an initialization for rapid task adaptation using only one or a few steps of gradient descent updates. In FSRL, each task is relation-centric, comprising only a few training triplets of a single relation, and the objective is to predict instances of novel relations not seen during meta-training. These methods focus on learning and encoding meta-knowledge across tasks to support rapid adaptation with minimal supervision.
Limitations of prior work. Despite promising results, current meta-learning based FSRL methods face two key limitations. First, they learn about meta-knowledge within individual tasks while neglecting shared patterns that exist across tasks. This limitation arises from the assumption that meta-training tasks are independent and identically distributed (i.i.d.), where a base model is adapted to each meta-training task using its support set, and then updated based on the performance on its query set. However, KG relations are heterogeneous and often semantically clustered, with some relations more closely related than others. For example, on Nell-One (Mitchell et al., 2018), FatherOfPerson and BrotherOf share a strong thematic overlap (family ties), whereas ColorOf pertains to a different type of attribute (physical properties). Ignoring such underlying commonalities among related tasks hinders model generalization in few-shot scenarios. While accounting for task variance has been explored in other few-shot domains like image classification (Li et al., 2019; Zhou et al., 2021), it remains largely unaddressed in the unique, non-i.i.d. context of relational learning.
Second, current methods are limited by their reliance on a single set of global parameters to encode meta-knowledge. With a shared initialization, a uniform adaptation process struggles to capture the unique characteristics of individual tasks from only a few support triplets. This is critical because relations and entities in KGs exhibit diverse interaction patterns (Wang et al., 2014; Ji et al., 2015), such as one-to-one, one-to-many, many-to-one, and many-to-many. An entity often reveals entirely different aspects depending on the relational context. For instance, the entity Elon Musk participates in the CeoOf relation (a professional role) and the FatherOfPerson relation (a family role). Current methods, constrained by a shared starting point, cannot effectively adapt to these locally distant contexts. Failing to modulate meta-knowledge for a specific task results in suboptimal adaptation.
Present work and contributions. To address these limitations, we propose MoEMeta, a novel meta-learning framework to address a critical, non-trivial challenge unaddressed by prior FSRL work: how to disentangle globally shared knowledge from task-specific contexts to achieve effective generalization and rapid adaptation under few-shot settings. MoEMeta provides a principled solution through two novel components that work in synergy.
  • Global knowledge generalization: To capture what to generalize, we introduce a mixture-of-experts (MoE) model to dynamically learn common relational patterns shared across tasks, referred to as relational prototypes, from a global pool of experts. The MoE is globally optimized, enabling the dynamic selection of experts that specialize in modeling the compositional nature of relations in KGs—an explicit and novel design for enhancing generalization.
  • Local context adaptation: To address what to adapt, we propose a novel task-tailored adaptation mechanism that leverages task-specific structural contexts to fine-tune relation-meta and entity embeddings for local adaptation. Unlike general task-aware modulation methods (Vuorio et al., 2019; Abdollahzadeh et al., 2021), which rely on task mode prediction to modulate the meta-learned prior parameters, our adaptation operates at the embedding level, offering a lightweight yet effective way to capture diverse entity-relation interactions unique to individual tasks.
Together, these two innovations achieve a good balance between global knowledge generalization and fine-grained task-specific adaptation, which is critical for FSRL with limited supervision. Experimental results demonstrate that MoEMeta achieves state-of-the-art performance on three KG benchmarks, significantly outperforming existing methods. In-depth ablation studies and visualizations further validate the efficacy of each component in MoEMeta.
2 Related Work
Few-Shot Relation Learning.
Existing approaches to FSRL can be categorized into metric-learning methods and meta-learning methods. Metric-learning methods learn a similarity metric between the support and query sets for adaptation. Approaches in this category include GMatching (Xiong et al., 2018), designed for one-shot relation learning; FSKGC (Zhang et al., 2020), which extends to few-shot settings using a recurrent autoencoder to aggregate support instances; FAAN (Sheng et al., 2020), which employs a dynamic attention mechanism for improved one-hop neighbor aggregation; CSR (Huang et al., 2022), which utilizes subgraph matching for adaptation; and NP-FKGC (Luo et al., 2023), which leverages neural processes to address distribution shifts in unseen relations.
Most mainstream FSRL methods are based on meta-learning (Finn et al., 2017), which learns an initial meta-prior from existing tasks for rapid adaptation to new few-shot relations. MetaR (Chen et al., 2019) learns relation-specific meta-information by simply averaging support triplet representations. MetaP (Jiang et al., 2021) utilizes meta-pattern learning under one-shot settings. GANA (Niu et al., 2021) integrates meta-learning with TransH (Wang et al., 2014) and devises a gated and attentive neighbor aggregator to address sparse neighbors. HiRe (Wu et al., 2023) exploits multi-granularity relational information to learn more generalizable meta-knowledge. RelAdapter (Ran et al., 2024) adopts a feedforward network with a residual layer to adapt relation-meta to target relation. Despite their progress, these methods face two key limitations: (1) they often neglect common relational patterns when learning meta-knowledge during meta-training, and (2) they fail to effectively incorporate local, task-specific contexts essential for rapid adaptation. Our work addresses these critical gaps by proposing a principled meta-learning framework that disentangles globally shared knowledge from task-specific adaptation, enabling both effective generalization and fast adaptation in FSRL. For a more extensive review of related work, including KG embedding learning and general few-shot learning techniques, please refer to Appendix C.
Mixture of Experts.
The mixture-of-experts (MoE) architecture, originally introduced by Jacobs et al. (1991), is a modular machine learning framework designed to improve performance by dynamically combining multiple specialized models, known as “experts”. Sparse MoEs and their “soft” variants (Jordan and Jacobs, 1994; Puigcerver et al., 2024) have been widely adopted in deep learning, especially in large-scale models for natural language processing and computer vision, to improve model training efficiency and scalability through adaptive subnetwork selection or dynamic pruning (Shazeer et al., 2017; Du et al., 2022). In a related KG context, MoE has been applied to fuse different semantics of relations (structural ontology and textual descriptions) at the triplet level (Song et al., 2022). However, our work innovatively proposes MoE as a meta-learner for FSRL, operating at the task level to learn specialized experts that can capture diverse relational patterns, thereby achieving superior task adaptation and knowledge transfer.
3 Preliminaries
3.1 FSRL Problem Formulation
A knowledge graph (KG) can be represented as a collection of factual triplets $\mathcal{G}=\{(h,r,t)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ denotes the set of entities, and $\mathcal{R}$ denotes the set of relations. Each triplet $(h,r,t)$ indicates a relationship $r$ between head entity $h$ and tail entity $t$.
Given a novel relation $r\notin\mathcal{R}$ and its corresponding support set $\mathcal{S}_{r}=\{(h_{i},r,t_{i})\}_{i=1}^{K}$, where $h_{i},t_{i}\in\mathcal{E}$. The objective of few-shot relational learning (FSRL) is to predict the missing tail entities for a query set $\mathcal{Q}_{r}=\{(h_{j},r,?)\}$, given a head entity $h_{j}\in\mathcal{E}$ and a relation $r$. A candidate set $\mathcal{C}_{h_{j},r}$ is provided for each query triplet $(h_{j},r,?)\in\mathcal{Q}_{r}$. The goal is to rank the true tail entity highest among all candidates. The support set $\mathcal{S}_{r}$ and query set $\mathcal{Q}_{r}$ jointly define a task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})$ for relation $r$. In $K$-shot relation learning, the support set $\mathcal{S}_{r}$ contains only a few training triplets ($K=1,3,5$).
3.2 Meta-Learning based FSRL
Given the strength of meta-learning in enabling rapid adaptation with minimal data, we build our approach on a meta-learning based framework, involving two stages: meta-training and meta-testing. The model is expected to learn a meta-prior from seen tasks in the meta-training stage, which is then effectively adapted to novel tasks in the meta-testing stage.
Meta-Training Stage. During meta-training, an FSRL model is trained on a set of relations $\mathcal{R}_{\text{tr}}$ with corresponding task data $\mathcal{D}_{\text{tr}}=\{\mathcal{T}_{r}\mid r\in\mathcal{R}_{\text{tr}}\}$. The model aims to learn a prior, which initializes a relation-meta learner for a new task. For each task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})\in\mathcal{D}_{\text{tr}}$, the learner extracts a task-specific relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ based on the support set $\mathcal{S}_{r}$, which is refined using a gradient meta $\mathbf{G}_{\mathcal{T}_{r}}$ computed from the support loss:
$\mathbf{R}_{\mathcal{T}_{r}}$ $\leftarrow\mathbf{R}_{\mathcal{T}_{r}}-\alpha\mathbf{G}_{\mathcal{T}_{r}},$ (1)
where $\alpha$ is the learning rate. The adapted relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ is then applied to the query set with the ground-truth tail entities, where the query loss is backpropagated to update both the prior and the entity embedding matrix.
Meta-Testing Stage. During meta-testing, the model is tested on novel relations $\mathcal{R}_{\text{te}}$ and their associated task data $\mathcal{D}_{\text{te}}=\{\mathcal{T}_{r}\mid r\in\mathcal{R}_{\text{te}}\}$, where $\mathcal{R}_{\text{tr}}\cap\mathcal{R}_{\text{te}}=\emptyset$. For each novel task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})\in\mathcal{D}_{\text{te}}$, the model utilizes the learned prior to initialize the relation-meta learner and compute an initial relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which is then adapted based on the support set $\mathcal{S}_{r}$ to obtain $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$. Finally, for each query triplet $(h,r,?)$, the model ranks all candidate tail entities $t$ in the candidate set $\mathcal{C}_{h,r}$ by computing a score $\textrm{s}(h,\mathbf{R}_{\mathcal{T}_{r}},t)$ for each candidate and sorting them in ascending order.
4 The Proposed MoEMeta Framework
This section presents our proposed MoEMeta framework. Its design is underpinned by two key insights: First, relations in real-world KGs exhibit heterogeneous characteristics, with some subsets of relations sharing greater similarity and stronger interconnections; Second, relations in KGs manifest diverse interaction patterns, such as one-to-many, many-to-one, and many-to-many. Building upon these insights, we introduce MoEMeta, a novel meta-learning framework that (1) utilizes an MoE model to dynamically learn relational prototypes and generate relation-meta for each task; (2) incorporates a task-tailored adaptation mechanism to facilitate local adaptation to each task.
Figure 1: The computation of the support loss. The support loss $\mathcal{L}(\mathcal{S}_{r})$ is used to update relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which is applied to the query set for model optimization.
MoEMeta consists of three core components: (1) attentive neighbor aggregation, (2) MoE-based meta-knowledge learning, and (3) task-tailored local adaptation. Figure 1 illustrates the detailed process of calculating the loss on the support set, through these three components. In essence, $\text{MoEMeta}(\bm{\Phi},\bm{\eta})$ contains two sets of learnable parameters: the parameters in attentive neighbor aggregation and MoE-based meta-knowledge learning are globally optimized during meta-training, denoted as $\bm{\Phi}$; while the parameters in task-tailored local adaptation are locally optimized within each task, denoted as $\bm{\eta}$.
4.1 Attentive Neighbor Aggregation
Given a $K$-shot task $\mathcal{T}_{r}$ with its support set $\mathcal{S}_{r}=\{(h_{i},r,t_{i})\}_{i=1}^{K}$, we propose an attentive neighbor encoder that aggregates the neighborhood information for both head entity $h_{i}$ and tail entity $t_{i}$. Specifically, for each entity, the encoder processes its neighboring tuples (neighboring entities and relations), combined with its own embeddings, to generate an enriched representation.
For a given entity $e$, its one-hop neighbors are represented as a set of neighboring tuples $(r_{i},e^{\prime}_{i})\}_{i=1}^{n}$, where $n$ is the total number of neighbors, $e^{\prime}_{i}$ is the neighboring entity and $r_{i}$ denotes the relation between $e^{\prime}_{i}$ and $e$. We first concatenate the embeddings of $r_{i}$ and $e^{\prime}_{i}$ to form the embedding $\mathbf{c}_{i}$ of the $i$-th tuple, i.e.  $\mathbf{c}_{i}=\mathbf{r}_{i}\oplus\mathbf{e}^{\prime}_{i}$, where $\oplus$ denotes concatenation of two vectors. To capture interactions among neighbors, we then apply a linear transformation followed by a non-linear activation function:
$\mathbf{c}^{\prime}_{i}=\textrm{ReLU}(\mathbf{W}\mathbf{c}_{i}),$ (2)
where $\mathbf{W}$ is a learnable weight matrix. To enable the model to attend to informative neighbors, we compute an attention-like gate value $g_{i}$ for each neighboring tuple using a sigmoid function:
$g_{i}=\sigma(\bm{\beta}^{T}\mathbf{c}^{\prime}_{i}),$ (3)
where $\bm{\beta}$ is a learnable parameter vector. The gate values control the influence of different neighbors in the aggregation process. Finally, the weighted neighbor embeddings are aggregated and combined with the entity’s own embedding $\mathbf{e}$ to obtain the final embedding:
$\mathbf{e}=\text{Aggregate}(\{g_{i}\cdot\mathbf{c}^{\prime}_{i}\}_{i=1}^{n})+\mathbf{e},$ (4)
The neighbor encoder is applied to both head entity $h$ and tail entity $t$ independently to ensure that the final embedding reflects both the entity’s intrinsic information and its local neighborhood structure.
4.2 MoE-based Meta-Knowledge Learning
To overcome the limitation of existing methods that overlook shared meta-knowledge across related tasks, we introduce a globally learned MoE model that captures common relational patterns, termed relational prototypes. For each task, a subset of relevant experts is dynamically selected from a shared pool based on its support triplets to learn relation-meta. This mechanism enables adaptive relation decomposition and expert specialization, facilitating the learning of diverse, compositional relational patterns and promoting effective generalization across tasks.
Specifically, we design a sparsely-gated MoE model consisting of $M$ experts. Each expert within the MoE is architecturally identical to a standard MLP (multilayer perceptron). For a given support triplet, a gating mechanism dynamically routes the triplet to a selected subset of experts, assigning weights to each expert in the subset based on its relevance. The outputs of the top-$N$ experts are then aggregated according to their weights, resulting in a representation of the relation associated with the triplet. To compute the relation-meta for a support set, we average the relation representations derived from all triplets within the support set. This approach enables the MoE layer to retrieve the most relevant meta-knowledge from the shared expert pool to model diverse relational patterns.
Given a $K$-shot task $\mathcal{T}_{r}$ and its support set $\mathcal{S}_{r}=\{(h_{i},r,t_{i})\}_{i=1}^{K}$, the head and tail entities of each triplet, $h_{i}$ and $t_{i}$, are augmented via our attentive neighbor aggregator, denoted as $\mathbf{h}_{i}$ and $\mathbf{t}_{i}$. The relation representation $\mathbf{r}_{i}$ derived from $(h_{i},t_{i})$ can be computed by our MoE as follows:
$\mathbf{r}_{j}$ $=\frac{1}{N}\sum_{i=1}^{M}g_{i,j}\text{MLP}_{i}(\mathbf{h}_{j},\mathbf{t}_{j};\bm{\theta}_{i}),$ (5)
$g_{i,j}$ $=\begin{cases}s_{i,j},&\text{if }s_{i,j}\in\text{Top}_{j}\left(\{s_{i,j}\mid 1\leq i\leq M\},N\right),\\ 0,&\text{otherwise},\end{cases}$ (6)
$s_{i,j}$ $=\text{softmax}(\text{Gate}(\mathbf{h}_{j},\mathbf{t}_{j};\bm{\theta}_{g})),$ (7)
where $M$ denotes the total number of experts, each implemented as a two-layer MLP, $\text{MLP}_{i}(\cdot;\bm{\theta}_{i})$. For each support triplet $j$, a gate network, $\text{Gate}(\cdot;\bm{\theta}_{g})$, which is also a two-layer MLP, computes a relevance score $s_{i,j}$ for each expert $i$. These scores are normalized via a softmax function over all $M$ experts. The function $\text{Top}_{j}(\cdot,N)$ selects only the top $N$ highest scores among all $M$ experts for each triplet $j$ and setting the rest to zero. This sparsity mechanism enables the model to selectively activate only a subset of experts for each support triplet, thereby facilitating the learning of more specialized relational prototypes globally shared across tasks.
The final relation-meta is computed as the average of relation representations from all support triplets:
$\mathbf{R}_{\mathcal{T}_{r}}=\frac{1}{K}\textstyle\sum_{i=1}^{K}\mathbf{r}_{i}$ (8)
4.3 Task-Tailored Local Adaptation
[…]
Specifically, for each task, MoEMeta maintains a task-specific set of local parameters, denoted as $\bm{\eta}$, including three randomly initialized projection vectors, $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$, which project the embeddings $\mathbf{h}_{j}$, $\mathbf{R}_{\mathcal{T}_{r}}$, and $\mathbf{t}_{j}$ into a task-specific space. Inspired by TransD (Ji et al., 2015), which addresses the limitations of fixed mapping strategies using per-entity and per-relation projections, our approach leverages task-specific projections to avoid overparametrization under few-shot settings. As all triplets within a task share the same relation, we adopt a unified projection vector $\mathbf{p}_{h}$ for all head entities, and $\mathbf{p}_{t}$ for all tail entities within the task. This provides a lightweight solution for enabling efficient task-specific adaptation with minimal parameter overhead.
Unlike previous FSRL methods including RelAdapter (Ran et al., 2024), which do not account for diverse interactions between entities and relations tailored to different tasks, our approach allows for a relation-specific adaptation, with projection operations defined below:
$\mathbf{h}_{j}^{\prime}$ $=\mathbf{h}_{j}+(\mathbf{p}_{h}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}},$ (9)
$\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$ $=\mathbf{R}_{\mathcal{T}_{r}}+(\mathbf{p}_{r}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}},$ (10)
$\mathbf{t}_{j}^{\prime}$ $=\mathbf{t}_{j}+(\mathbf{p}_{t}^{\top}\mathbf{R}_{\mathcal{T}_{r}})\cdot\mathbf{R}_{\mathcal{T}_{r}}.$ (11)
Above, the dot product $\mathbf{p}^{\top}\mathbf{R}_{\mathcal{T}_{r}}$ yields a scalar alignment score between each projection vector $\mathbf{p}$ and relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which then scales $\mathbf{R}_{\mathcal{T}_{r}}$ to generate an offset vector that modulates specific features in the entity embeddings and relation-meta. Although the three projection vectors can functionally be merged into a single perturbation term, maintaining separate projections for head entities, relation, and tail entities enhances modularity and flexibility, allowing for more expressive or heterogeneous projection functions that can incorporate richer inductive biases or task-specific modeling needs in future extensions. The updated embeddings ($\mathbf{h}_{j}^{\prime}$, $\mathbf{t}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$) thus capture task-specific nuances while maintaining consistency with the generalizable meta-knowledge. This local adaptation scheme ensures that each task can adaptively focus on its unique aspects but also preserve the shared commonalities across related tasks.
To further optimize the adapted embeddings $\mathbf{h}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{j}^{\prime}$, we define a scoring function in the projected space to measure the plausibility of each triplet $(h_{j},r,t_{j})$, which is given by:
$\mathrm{score}(h_{j},t_{j})=\left\lVert\mathbf{h}_{j}^{\prime}+\mathbf{R}_{\mathcal{T}_{r}}^{\prime}-\mathbf{t}_{j}^{\prime}\right\rVert_{2},$ (12)
where $\lVert\cdot\rVert_{2}$ denotes the $\ell_{2}$ norm. $\mathbf{h}_{j}^{\prime}$, $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$, and $\mathbf{t}_{j}^{\prime}$ are the projected embeddings of the head entity, relation, and tail entity, respectively. This formulation ensures that the distance between the adapted head entity embedding $\mathbf{h}_{j}^{\prime}$ and the adapted tail entity embedding $\mathbf{t}_{j}^{\prime}$ is minimized after being translated by the relation embedding $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$ in the task-specific projection space.
Considering the entire support set, we define a margin-based loss function as follows:
$\mathcal{L}(\mathcal{S}_{r})=\sum_{(h_{j},r,t_{j})\in\mathcal{S}_{r}}\max\{0,\mathrm{score}(h_{j},t_{j})+\gamma-\mathrm{score}(h_{j},t^{\prime}_{j})\},$ (13)
where $\gamma$ is a margin hyperparameter, and $(h_{j},r,t^{\prime}_{j})$ represents a negative triplet generated by replacing the tail entity $t_{j}$ with a random entity $t^{\prime}_{j}$ such that $(h_{j},r,t^{\prime}_{j})\notin\mathcal{G}$.
Once the gradients of the loss function $\mathcal{L}(\mathcal{S}_{r})$ are computed, we optimize the projection vectors $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, $\mathbf{p}_{t}$, together with the relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, to ensure generalization to the query set. Specifically, the projection vectors can be updated in a unified manner as follows:
$\mathbf{p}_{x}$ $\leftarrow\mathbf{p}_{x}-\alpha\frac{\partial\mathcal{L}(\mathcal{S}_{r})}{\partial\mathbf{p}_{x}},\quad x\in\{h,r,t\}.$ (14)
Similarly, the relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ is updated as:
$\mathbf{R}_{\mathcal{T}_{r}}$ $\leftarrow\mathbf{R}_{\mathcal{T}_{r}}-\alpha\frac{\partial\mathcal{L}(\mathcal{S}_{r})}{\partial\mathbf{R}_{\mathcal{T}_{r}}},$ (15)
where $\alpha$ is the learning rate. These updates adapt the model based on the support set, which is then evaluated on the query set $\mathcal{Q}_{r}$. For each query triplet $(h_{q},r,t_{q})$, we evaluate its plausibility for each $t_{q}$ from the candidate set $C_{h_{q},r}$, using the score function Eq. (12). Then, the query loss is defined as:
$\mathcal{L}(\mathcal{Q}_{r})=\sum_{(h_{q},r,t_{q})\in\mathcal{Q}_{r}}\max\bigl\{0,\mathrm{score}(h_{q},t_{q})+\gamma-\mathrm{score}(h_{q},t^{\prime}_{q})\bigr\},$ (16)
where $(h_{q},r,t^{\prime}_{q})$ is a negative triplet generated by corrupting the tail entity $t_{q}$ with a randomly sampled entity $t^{\prime}_{q}$. The optimization objective for training the entire model is to minimize $\mathcal{L}(\mathcal{Q}_{r})$.
The task-specific projection vectors $\mathbf{p}_{h}$, $\mathbf{p}_{r}$, and $\mathbf{p}_{t}$ are randomly initialized and trained specifically for each task on the support set $\mathcal{S}_{r}$, enabling the model to adapt locally to each task. This mechanism ensures that (1) task-specific adaptation: By projecting embeddings into a task-specific latent space, the model captures fine-grained task-specific local contexts. (2) complementarity to MoE: While the MoE captures globally shared relational prototypes, dynamic learnable projections provide finer-grained local adaptation, thereby improving model generalization under few-shot settings.
4.4 MoEMeta Meta-Training and Meta-Testing
During meta-training, MoEMeta optimizes two sets of parameters: global parameters $\bm{\Phi}=\{\mathbf{W},\bm{\beta},\{\bm{\theta}\}_{i=1}^{M},\bm{\theta}_{g}\}$, for attentive neighbor aggregation and MoE-based meta-knowledge learning, which are globally optimized; and task-specific adaptation parameters $\bm{\eta}=\{\mathbf{p}_{h},\mathbf{p}_{r},\mathbf{p}_{t}\}$, which are optimized locally for each task. The meta-training algorithm of MoEMeta is detailed in Algorithm 1.
During meta-testing, the learned global parameters $\bm{\Phi}$ are used to initialize the model and remain frozen. For each novel task, only the local parameters $\bm{\eta}$ are fine-tuned on the support set before evaluation on the query set by ranking candidate tail entities, following the same inference procedure as in meta-training. The complexity analysis of MoEMeta’s components are provided in Appendix A.
Input: Knowledge Graph $\mathcal{G}$
Output: Optimized model parameters $\bm{\Phi}$
1
2Initialize globally optimized parameters $\bm{\Phi}=\{\textbf{W},\bm{\beta},\{\bm{\theta}\}_{i=1}^{M},\bm{\theta}_{g}\}$;
3
4while not converged do
5   Sample $\mathcal{T}_{r}$ = $\mathcal{S}_{r}$, $\mathcal{Q}_{r}$};
6 foreach triplet $(h_{i},r,t_{i})\in\mathcal{S}_{r}$ do
    // Attentive Neighbor Aggregation
7      Attentively aggregate neighbors using Eq. 2 -Eq. 4;
    // MoE-based Meta-Knowledge Learning
8      Compute expert selection scores using Eq. 7;
9      Select top-NN experts and compute $\mathbf{r}_{i}$;
10
11   end foreach
12  Compute relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ using Eq. 8;
13
 // Task-Tailored Local Adaptation
14   Initialize locally optimized parameters $\bm{\eta}=\{\mathbf{p}_{h},\mathbf{p}_{r},\mathbf{p}_{t}\}$;
15 foreach triplet $(h_{i},r,t_{i})\in\mathcal{S}_{r}$ do
16      Compute projected embeddings using Eq. 9–Eq. 11;
17
18   end foreach
 // Inner loop optimization
19   Compute the support loss $\mathcal{L}(\mathcal{S}_{r})$ using Eq. 13;
20   Optimize projection vectors using Eq. 14;
21   Optimize relation-meta representation using Eq. 15;
22
 // Evaluation on Query Set
23   Compute the query loss for all $(h_{q},r,t_{q})\in\mathcal{Q}_{r}$ via Eq. 16;
 // Optimization
24   Optimize model parameters $\bm{\Phi}$ based on the query loss.
25 end while
Algorithm 1 Meta-training Algorithm of MoEMeta
5 Experiments
Datasets. We evaluate our method on three widely used KG benchmarks for few-shot relational learning: Nell-One and Wiki-One (Xiong et al., 2018), as well as FB15K-One (Ran et al., 2024). The training/validation/test splits include 51/5/11, 133/16/34, and 75/11/33 tasks on Nell-One, Wiki-One, and FB15K-One, respectively. Detailed dataset statistics are summarized in Table 1.
Table 1: Statistics of datasets.
Dataset # Relations # Entities # Triplets # Tasks
Nell-One 358 68,545 181,109 67
Wiki-One 822 4,838,244 5,859,240 183
FB15K-One 237 14,541 281,624 119
Evaluation Metrics. We report both MRR (mean reciprocal rank) and Hits@k ($k=1,5,10$). MRR captures the average reciprocal rank of correct entities, while Hits@k measures the proportion of correct entities appearing in the top $k$ predictions. Following prior studies, our method is evaluated against baselines under 1-shot and 5-shot settings on Nell-One and Wiki-One, which are commonly used for evaluating FSRL, and under the 3-shot setting on FB15K-One, following the setup of RelAdapter (Ran et al., 2024).
Table 2: Comparison against state-of-the-art methods on Nell-One and Wiki-One. MetaR-I and MetaR-P indicate the In-train and Pre-train of MetaR (Chen et al., 2019).
[…]
Baselines. We compare our proposed method with two categories of state-of-the-art baselines:
  • Conventional KG embedding methods: This group focuses on modeling relational structures in KGs to learn entity and relation embeddings. We compare with the widely-used baselines including TransE (Bordes et al., 2013), TransD (Ji et al., 2015), DistMult (Yang et al., 2015), and ComplEx (Trouillon et al., 2016). Using OpenKE (Han et al., 2018), these models are reproduced with hyperparameters specified in the original papers.
  • Few-shot relational learning methods: We compare our method against state-of-the-art FSRL baselines evaluated under the same framework, (1) metric learning methods: GMatching (Xiong et al., 2018), FAAN (Sheng et al., 2020), FSKGC (Zhang et al., 2020); and (2) strong meta-learning competitors: MetaR (Chen et al., 2019), GANA (Niu et al., 2021), HiRe (Wu et al., 2023), and RelAdapter (Ran et al., 2024). We exclude methods like CSR (Huang et al., 2022) and NP-FKGC (Luo et al., 2023) as they employ a different training and evaluation protocol.
All results are obtained after the models are trained using all triplets from background relations (Xiong et al., 2018) and meta-training tasks, and evaluated on meta-testing tasks.
Experimental Setup. Following prior work, we set the embedding dimension to $100$ for Nell-One and FB15K-One, and $50$ for Wiki-One, initialized using TransE-pretrained weights provided by GMatching (Xiong et al., 2018). The number of neighbors per entity is capped at 50 for aggregation. For MoE, each expert is a two-layer MLP (hidden dimension: 64, ReLU). The gate network is also a two-layer MLP (hidden dimension: 64, output dimension: 1, and ReLU). We set the number of experts $M$ as $32$ and the number of selected experts $N$ as $5$. The margin $\gamma$ in Eq. 16 is set to $1$. MoEMeta is trained using Adam (batch size: $1,024$, learning rate: $0.001$). All results are averaged over 5 independent runs with different random seeds. All models are implemented in PyTorch and trained on a single Tesla P100 GPU. More reproducibility details are provided in Appendix B.1.
5.1 Comparison with Baseline Methods
Table 3: Comparison under 3-shot setting on FB15K-One.
FB15K-One
Methods MRR Hits@1 Hits@10
TransE 0.294 0.204 0.437
TransD 0.301 0.203 0.439
DistMult 0.234 0.208 0.364
ComplEx 0.239 0.205 0.359
GMatching 0.309 0.245 0.441
MetaR 0.368 0.251 0.536
FAAN 0.363 0.279 0.542
GANA 0.388 $\underline{0.301}$ 0.553
HiRe 0.378 0.281 0.571
RelAdapter $\underline{0.405}$ 0.297 $\underline{0.575}$
MoEMeta 0.423 0.302 0.651
Tables 2 and 3 compare MoEMeta against state-of-the-art methods on Nell-One and Wiki-One under 1-shot and 5-shot settings, and on FB15K-One under 3-shot setting. The best and the second best performers are respectively marked by grey and underline. Our proposed MoEMeta consistently achieves state-of-the-art results across all datasets and settings. On Nell-One, MoEMeta surpasses HiRe, the second best performer, by 11.8% MRR, 23.9% Hits@1, 6.4% Hits@5, and 6.1% Hits@10 under $1$-shot setting, and 10.8% MRR, 14.0% Hits@1, 2.3% Hits@5, and 0.6% Hits@10 under $5$-shot setting. On FB15K-One, MoEMeta outperforms the second best method by 4.4% MRR, 0.3% Hits@1, and 13.2% Hits@5 under $3$-shot setting. We further conduct paired t-tests between MoEMeta and the second best performers on all datasets. The improvements of MoEMeta over the second best method with statistical significance at the 0.05 level are marked with . These results highlight MoEMeta’s effectiveness in learning generalizable meta-knowledge that adapts well to novel relations and proving its competitiveness against other baselines.
Across three datasets, the most substantial improvements are on Nell-One, particularly in MRR under both 1-shot and 5-shot settings. The primary gain is attributed to the increase in the challenging Hits@1 metric, which gauges the model’s ability to rank the correct tail entity as the top candidate among all choices. This is especially notable since triplets in Nell-One are constructed in natural language, providing richer semantics that can be more effectively captured by our relational prototypes. These results validate the efficacy of our proposed MoE-based relation-meta learner in learning globally shared, fine-grained relational patterns and adapting these prototypes to new, unseen tasks. We additionally provide a runtime analysis for both meta-training and meta-testing in Appendix B.2.
5.2 Ablation Studies
Table 4: Ablation study on Nell-One and FB15K-One.
                        Nell-One FB15K-One Ablation on 1-shot 5-shot 3-shot MRR Hits@1 Hits@5 Hits@10 MRR Hits@1 Hits@5 Hits@10 MRR Hits@1 Hits@10 MoEMeta 0.322 0.228 0.429 0.501 0.339 0.236 0.440 0.523 0.422 0.298 0.652 w/o N.A 0.311 0.219 0.418 0.489 0.328 0.231 0.436 0.509 0.413 0.286 0.631 w/o MoE 0.291 0.199 0.392 0.467 0.293 0.201 0.399 0.484 0.401 0.273 0.624 w/o L.A 0.301 0.206 0.398 0.476 0.315 0.213 0.415 0.501 0.408 0.285 0.628 N.A indicates “attentive neighbor aggregation” and L.A indicates “task-tailored local adaptation”.
To assess the contribution of MoEMeta’s three key components, attentive neighbor aggregation, mixture of experts (MoE), and task-tailored local adaptation, we report the ablation study of MoEMeta on Nell-One under 1-shot and 5-shot settings, and FB15K-One under 3-shot setting in Table 4.
  • Removing Neighbor Aggregation (w/o N.A), which removes the first term in Eq. 4, results in a moderate performance decrease across all settings. This suggests the importance of incorporating neighboring information to enhance entity representations.
  • Replacing MoE (w/o MoE) with an MLP-based relation-meta learner from (Chen et al., 2019) leads to a significant performance decline; MRR drops by 9.6% and 13.6% under 1-shot and 5-shot settings on Nell-One, and 5.0% under 3-shot setting on FB15K-One. This underscores MoE’s indispensable role in acquiring relational prototypes as generalizable meta-knowledge.
  • Ablating task-tailored Local Adaptation (w/o L.A), where the learned relation-meta is directly applied to the query set, also causes a significant performance drop across datasets. This suggests that global meta-knowledge alone is insufficient, highlighting the crucial role of task-tailored adaptation for capturing local distinct patterns and improving generalization across tasks.
Figure 2: Two gating value distributions over 32 experts generated by our MoE-based meta-knowledge learner. The upper figure shows that three similar relations exhibit similar gating value distributions, suggesting that a similar set of experts are activated. In contrast, the lower figure demonstrates that for relations with distinct relational patterns, the activated experts differ significantly. This highlights the presence of shared relational patterns, underscoring the necessity of learning relation prototypes.
Figure 3: Performance analysis under 5-shot setting on Nell-One for1-N, N-1 and N-N relation types.
Methods Rel. types MRR Hits@1 Hits@5 Hits@10
MoEMeta 1-N 0.406 0.310 0.506 0.577
N-1 0.945 0.913 0.986 0.986
N-N 0.404 0.279 0.536 0.618
w/o L.A 1-N 0.391 0.292 0.500 0.565
N-1 0.874 0.871 0.986 0.986
N-N 0.372 0.256 0.485 0.586
Figure 4: Hyperparameter sensitivity study of MoE under 5-shot setting on Nell-One.
5.3 Analysis on Complex Relations
To further evaluate the significance of task-tailored local adaptation for complex relations, we conduct a detailed analysis on Nell-One under the 5-shot setting by categorizing meta-testing set into three distinct relation types: 1-N, N-1, and N-N. We compare MoEMeta with a variant without task-tailored local adaptation (w/o L.A). As shown in Table 4, the removal of local adaptation leads to a substantially larger performance drop on 1-N and N-N relations than on N-1 relations. This is because the tail entity in N-1 relations is relatively predictable, reducing the search space and enabling the model to learn a more stable mapping pattern. In contrast, 1-N and N-N relations exhibit greater ambiguity and a more dispersed distribution of tail entities, making them harder to predict. Notably, removing task-tailored adaptation results in the most significant performance drop in the N-N relation type. This analysis confirms the critical role of our task-tailored adaptation mechanism in enhancing generalization to complex relational patterns.
For better elaboration, we further visualize the distribution of gating values of different relations in Figure 2 to provide a more intuitive understanding of why learning relational prototypes would work. As observed, relations with similar patterns exhibit a comparable gating value distribution and activate a similar subset of experts, whereas distinct relations demonstrate an opposite pattern, highlighting the model’s ability to differentiate between diverse relational patterns but capture similar relational prototypes that are globally shared across related tasks.
5.4 Hyperparameter Analysis on MoE
To assess their impact of two MoE hyperparameters, the number of experts $M$ and the number of top selected experts $N$, we conduct a sensitivity analysis on Nell-One under 5-shot settings as a case study. When fixing $N=5$ and varying $M\in\{16,32,64\}$, MoEMeta achieves the best performance with 32 experts. This aligns with the guideline of setting the number of experts $M$ smaller than the total number of tasks to encourage learning shared relational prototypes and prevent overfitting with limited support triplets. When fixing $M=32$ and varying $N\in\{3,5,10\}$, we observe that selecting slightly more experts (e.g., 5 vs. 3) leads to improved performance, yet increasing $N$ to 10 causes only a slight drop, suggesting that MoEMeta is relatively insensitive to $N$ within a reasonable range.
6 Conclusion
This paper presents MoEMeta, a novel meta-learning framework for few-shot relational learning in KGs. By integrating an MoE model to capture relational prototypes and a task-tailored adaptation mechanism to refine embeddings locally, MoEMeta significantly enhances the model’s ability to generalize across tasks while facilitating rapid adaptation tailored to individual tasks. Experimental results on three real-world KG datasets demonstrate that MoEMeta consistently outperforms state-of-the-art methods, showcasing its effectiveness for few-shot relation learning. In-depth ablation studies and visualizations further validate the importance of MoEMeta’s key components, confirming its efficacy in handling complex relational data in KGs with limited supervision.
Appendix A Complexity Analysis
We analyze the computational complexity of our proposed MoEMeta for its key components. Table 5 presents the computational complexity of all components of MoEMeta, where $d$ denotes the dimension of entity embeddings, $n$ denotes the number of neighbors in attentive neighbor aggregation, and $K$ denotes the number of support triplets per task. For MoE, $M$ is the total number of experts, while $N$ is the number of activated experts in every forward pass.
As shown in Table 5, the complexity of MoEMeta scales linearly with the number of neighbors during the aggregation stage ($\mathcal{O}(Knd^{2})$) and linearly with the number of support triplets during MoE-based meta-knowledge learning ($\mathcal{O}(NKd^{2})$). The task-tailored local adaptation introduces only a minimal overhead of ($\mathcal{O}(d)$) in both parameters and computation. Overall, the computational cost is dominated by neighbor aggregation ($\mathcal{O}(Knd^{2})$), which is an essential component in current state-of-the-art methods. The additional computational cost and parameter overhead introduced by MoEMeta are negligible, making it efficient for few-shot KG relational learning.
Table 5: Complexity analysis of MoEMeta components with respect to the number of model parameters and the number of multiplication operations in each epoch. $d$ denotes the dimension of entity embeddings. $n$ denotes the number of neighbors per entity, and $K$ denotes the number of support triplets per task, where $K\ll n,d$. $M$ is the total number of experts, while $N$ is the number of activated experts in every forward pass.
Component # Parameters # Multiplication Operations
Attentive neighbor aggregation $\mathcal{O}(d^{2})$ $\mathcal{O}(Knd^{2})$
MoE-based meta-knowledge learning $\mathcal{O}(Md^{2})$ $\mathcal{O}(NKd^{2})$
Task-tailored local adaptation $\mathcal{O}(d)$ $\mathcal{O}(d)$
Total $\mathcal{O}(Md^{2})$ $\mathcal{O}(nKd^{2})$
Appendix B Additional Experimental Details and Results
B.1 Reproducibility Details
Our method and all baselines are compared using the same evaluation protocol. The results of baseline methods are reproduced by using their official open-sourced implementations or adopted from their respective papers.
For MetaR (Chen et al., 2019) (both In-Train and Pre-Train) [^2] MetaR: https://github.com/AnselCmy/MetaR,FAAN (Sheng et al., 2020) [^3] FAAN: https://github.com/JiaweiSheng/FAAN, MetaP (Jiang et al., 2021) [^4] MetaP:https://github.com/jzystc/MetaP and HiRe (Wu et al., 2023) [^5] HiRe: https://github.com/alexhw15/HiRe, we directly use the results reported in their papers. The results for GMatching (Xiong et al., 2018) [^6] GMatching: https://github.com/xwhan/One-shot-Relational-Learning under 1-shot and 5-shot settings are taken from (Chen et al., 2019). Since FSKGC (Zhang et al., 2020) [^7] FSKGC: https://github.com/chuxuzhang/AAAI2020_FSRL was originally evaluated under different conditions (with a smaller candidate set), we use the re-implemented results provided by FAAN to ensure consistency. For GANA (Niu et al., 2021) [^8] GANA: https://github.com/ngl567/GANA-FewShotKGC, we reproduce its results by setting the number of neighbors to $50$, aligning with the settings of other methods. We reproduce the results of RelAdapter (Ran et al., 2024) [^9] RelAdapter: https://github.com/smufang/RelAdapter under the 1-shot and 5-shot settings on Wiki-One. However, as the pre-trained contextual information required for Nell-One is not released, the performance of RelAdapter on this dataset is unavailable for evaluation. Following RelAdapter, the comparison results on FB15K-One under 3-shot setting are reported in Table 3, with all baseline results sourced from RelAdapter (Ran et al., 2024). All models are implemented in PyTorch and trained on a single Tesla P100 GPU.
B.2 Runtime Analysis
We conduct a runtime analysis comparing our proposed MoEMeta with other baselines. For fair comparison, we evaluate MoEMeta on a RTX3090 GPU to keep consistency with baseline runtimes sourced from the RelAdapter (Ran et al., 2024) paper. The total runtime for meta-training and the inference time per triplet during meta-testing are reported in seconds in Table 6.
During meta-training, MoEMeta is only slightly slower than MetaR and RelAdapter, while being faster than other baselines such as GANA and FSKGC. MoEMeta requires 26,277 seconds to train on the NELL-One dataset, compared to 22,691 seconds for MetaR and 24,085 seconds for RelAdapter. The moderate increase in training time mainly comes from the two additional MLPs used in the MoE. Despite this additional overhead, MoEMeta achieves significantly better performance. For example, under the 1-shot setting on NELL-One, MoEMeta outperforms MetaR, by 28.8% MRR, 34.1% Hits@1, 27.7% Hits@5, and 24.9% Hits@10, respectively (see Table 2). For the inference during meta-testing, all methods including MoEMeta complete in milliseconds per triplet.
Table 6: Runtime comparison of meta-training and meta-testing time on Wiki-One and FB15K-One.
Methods Meta-training Time (seconds) Meta-testing Time per Triplet (seconds)
Wiki-One FB15K-One Wiki-One FB15K-One
GMatching 28,941 19,678 0.009 0.004
FSKGC 28,742 20,765 0.013 0.004
GANA 35,374 28,167 0.017 0.006
FAAN 32,036 23,675 0.016 0.005
HiRe 34,257 27,736 0.036 0.007
MetaR 22,691 16,504 0.012 0.005
RelAdapter 24,085 17,529 0.045 0.008
MoEMeta 26,277 18,321 0.019 0.006
Appendix C Detailed Related Work
C.1 Supervised KG Embedding Learning
KG embedding methods aim to learn a low-dimensional embedding space for entities and relations in KGs using input triplets as supervision. Two main categories of KG embedding methods for KG relational learning include: (1) Translation-based methods, such as TransE (Bordes et al., 2013), TransH (Wang et al., 2014), TransD (Ji et al., 2015), and TransR (Lin et al., 2015), which assumes a translational relationship between entities and relations for learning a unified embedding space. (2) Semantic matching methods, such as ComplEx (Trouillon et al., 2016), RotatE (Sun et al., 2019), and ComplEx-N3 (Lacroix et al., 2018), which improve the modeling of relation patterns in a vector/complex space. However, these methods rely on a large amount of triplets for training. Their performance deteriorates at few-shot settings, where only a handful of training triplets are available for each relation.
C.2 Few-Shot Learning and Meta Learning
Few-shot learning aims to train a supervised model that can generalize to new tasks with only limited amounts of training data (Fei-Fei et al., 2006). Meta-learning (Finn et al., 2017; Santoro et al., 2016; Ravi and Larochelle, 2017; Nichol and Schulman, 2018; Ren et al., 2018) has emerged as a powerful paradigm for enabling rapid learning on few-shot tasks by leveraging meta-knowledge learned from meta-training tasks. Relative to metric-based approaches (Jake Snell and Zemel, 2017; Vinyals et al., 2016), optimization-based meta-learning methods have gained significant attention due to their strong generalization capacities. These methods cast meta-learning into a bi-level optimization problem; the inner loop fine-tunes a base model on individual tasks, and the outer loop optimizes the base model for quick adaptation across tasks. MAML (Finn et al., 2017) is the most representative algorithm in this category, which learns global parameters as an initialization to be shared among task-specific models for fast adaptation within one or a few steps of gradient descent. Due to its flexible formulation, MAML has been applied to various domains including vision, robotics, lifelong learning, and beyond. Subsequent variants of MAML (Finn et al., 2018; Von Oswald et al., 2021) improve upon its adaptability by optimizing model initialization or learning adaptive learning rates for fast convergence (Sun and Gao, 2023; Zou et al., 2021; Finn et al., 2018). Our work focuses on learning well-generalized meta-knowledge across tasks while incorporating local adaptation to improve few-shot relational learning in KGs.
C.3 Few-Shot Relation Learning
Few-shot relation learning (FSRL) addresses the challenges of learning with limited training examples and relational information in KGs. Current FSRL approaches can be broadly categorized into two groups: metric-learning methods and meta-learning methods. Metric-learning methods focus on learning a similarity metric between the support and query sets for adaptation. GMatching (Xiong et al., 2018) introduces one-shot relation learning, where a one-hop neighbor encoder with equal weights is used to learn entity embeddings and a matching network is designed to compare similarity between support and query entity pairs. FSKGC (Zhang et al., 2020) generalizes the setting to more shots and employs a recurrent autoencoder to aggregate few-shot instances in the support set, yet imposing an unrealistic sequential dependencies among support triplets. FAAN (Sheng et al., 2020) proposes a dynamic attention mechanism to improve the aggregation of one-hop neighbors. CSR (Huang et al., 2022) employs subgraph matching as a hypothesis testing to adapt from the support triplets to the query set. NP-FKGC (Luo et al., 2023) employs neural process to better handle distribution shifts for unseen relations.
Most mainstream FSRL methods are based on meta-learning (Finn et al., 2017), which focuses on learning an initial meta-prior from existing tasks that can generalize to new few-shot relations. MetaR (Chen et al., 2019) learns relation-specific meta-information for adaptation, by simply averaging the representations of all support triplets. GANA (Niu et al., 2021) integrates meta-learning with TransH (Wang et al., 2014) and devises a gated and attentive neighbor aggregator to address sparse neighbors for learning informative entity embeddings. HiRe (Wu et al., 2023) exploits multi-granularity relational information to learn meta-knowledge that better generalizes to unseen relations. RelAdapter (Ran et al., 2024) adopts a feedforward network with a residual layer to adapt relation-meta to the target relation. Additionally, pre-trained context information of related entities are provided to augment entity embeddings.
Despite great progress, the aforementioned meta-learning based methods suffer from two key limitations. First, these methods neglect common relational patterns among diverse relations when learning meta-knowledge during meta-training. Second, they fail to effectively incorporate local, task-specific information essential for fast adaptation, as global meta-knowledge may not be optimally tailored to the unique characteristics of individual tasks. Our work is proposed to address these critical gaps within the meta-learning framework, enabling more effective model generalization and adaptation under few-shot settings.
Our work is also related to recent graph-based and contrastive meta-learning methods such as (Liu et al., 2022) and (Liu et al., 2024). While these approaches are relevant to few-shot learning and meta-learning on node-level tasks for homogeneous graphs, they are not directly applicable to our setting, which focuses on few-shot relational learning in KGs—that centers on relation-level tasks. In contrast, our method, MoEMeta, specifically targets the key challenges of few-shot relational learning, such as the complex and diverse relational patterns in KGs. This distinction sets our work apart from general graph few-shot learning or meta-learning methods that target node-level tasks.